OpenStudy (anonymous):

Solve for x: log5 x+log5 (x-4)=1. 5 is the base.

OpenStudy (loser66):

log a +log b = log (a*b) and 1 = log (base 5) of 5 then take off log both sides

OpenStudy (anonymous):

So x^2-4x=1, then you solve for x?

OpenStudy (dan815):